A tap supplies water at 26 degree Celsius while another supplies water at 82 degree. If the man wishes to bath with water at 40 degrees, find the ratio of the mass of hot water to that of cold water


1
Expert's answer
2021-02-17T11:01:37-0500
"M_1C(82 - 40) = M_2C(40 - 26)\\\frac{M_2}{M_1}=\frac{14}{42}=\frac{1}{3}"